OpenStudy (anonymous):

Think I got the answer help please. Find the average rate of change for the function over the given interval. y= x^2 between x=-5 and x=10 I think the answer may be -5

OpenStudy (johnweldon1993):

Well the average rate of change....is delta y / delta x delta x is the difference between your x-values 10 - (-5) = 15 is your delta x to find your delta y...you plug in both those numbers into the equation...and take the difference between the 2 values 10^2 = 100 -5^2 = 25 100 - 25 = 75 this is your delta y so delta y / delta x = 75 --- 15 which equals 5

OpenStudy (anonymous):

Thank-you

OpenStudy (johnweldon1993):

remember....delta y ...is y2 - y1 delta x is x2 - x1 x2 = 10 x1 = -5 y2 = 100 y1 = 25 so you would do y2 - y1 ------- x2 - x1 which came out to 5

OpenStudy (johnweldon1993):

and your welcome!